OpenStudy (anonymous):

tiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtired

OpenStudy (whpalmer4):

When you are confronted with a square root in the denominator with only one term, multiply top and bottom by the denominator and you'll eliminate it. When you have a binomial with a square root in the denominator, multiply top and bottom by the conjugate of the denominator. If denominator is (a+b), multiply by (a-b), and vice versa.

OpenStudy (anonymous):

you must multliply by the denominator's conjugate. and notice that technically I'm just multiplying by one. \[\frac{ 3 }{6 - \sqrt{5}} \times \frac{ 6 + \sqrt{5} }{ 6 + \sqrt{5} }\] just multiply and simplify for the final answer let me know if you'd like the next step ^_^

OpenStudy (anonymous):

that is the most simplified version. however, look at your denominator again. it is not 41 ^_^